# RAVEN: Rethinking Adversarial Video Generation with Efficient Tri-plane Networks

### v1

@@ Abstract
We present a novel unconditional video generative model designed to address long-term spatial and temporal dependencies. To capture these dependencies, our approach incorporates a hybrid explicit-implicit tri-plane representation inspired by 3D-aware generative frameworks developed for three-dimensional object representation and employs a singular latent code to model an entire video sequence. Individual video frames are then synthesized from an intermediate tri-plane representation, which itself is derived from the primary latent code. This novel strategy reduces computational complexity by a factor of $2$ as measured in FLOPs. Consequently, our approach facilitates the efficient and temporally coherent generation of videos. Moreover, our joint frame modeling approach, in contrast to autoregressive methods, mitigates the generation of visual artifacts. We further enhance the model’s capabilities by integrating an optical flow-based module within our Generative Adversarial Network (GAN) based generator architecture, thereby compensating for the constraints imposed by a smaller generator size. As a result, our model is capable of synthesizing high-fidelity video clips at a resolution of $256\times 256$ pixels, with durations extending to more than $5$ seconds at a frame rate of 30 fps. The efficacy and versatility of our approach are empirically validated through qualitative and quantitative assessments across three different datasets comprising both synthetic and real video clips.
1 Equal contribution.
@@ 1 Introduction
Figure 1: RAVEN can generate long-term spatially and temporally coherent videos. For visualization purposes, we only show $6$ equally spaced frames from a video clip of length $5.3$ seconds here.
The field of video generation has made remarkable progress in recent years, with the aim of generating high-quality content-guided videos that are both realistic and efficient. This progress, however, has not reached the same heights as image generation due to two primary challenges: the inherent complexity of video data and the intensive computational demand video processing entails.
In the realm of unconditional video generation, most existing approaches leverage high-performing image generation models, creating videos in a sequential, autoregressive manner within the latent space. This technique, while innovative, is widely accepted to often result in accumulation of error over the regression chain, often known as exposure bias [21, 7, 19, 1]. Image generators often are imperfect, even when they are carefully trained with underlying 3D structure in narrow domain [12]. This further complicates building autoregressive models that use pretrained image generators. Content-guided video generation, on the other hand, relies on additional modalities and is dependent on training with extensive datasets and complex models [37, 8].
@@
Our methodology begins by drawing parallels between the representation of 3D objects in neural networks and the challenges faced by video generative models. A 3D object, when modeled as a Neural Radiance Field (NeRF) [22], is described as a continuous volumetric scene function derived from a sparse set of images taken from different viewpoints. This volumetric representation is typically encapsulated by a parametric function that ingests a 5D coordinate—three for spatial location and two for viewing directions—and outputs the volume’s material density and view-dependent emitted radiance.
Extensive research has been conducted on parametric functions that represent location-dependent properties in three dimensions. By analogy, a video is a 3D array indexed by two spatial coordinates and one temporal coordinate. Thus, representing a video with a parametric function involves mapping a 3D input (x, y, t) to RGB pixel values, effectively parameterizing a video as a continuous volume of color.
When designing a generative model with a continuous output space, various forms can be employed, such as explicit voxel grids [9, 13] or implicit representations [5, 24], which are prevalent in 3D object modeling. However, these representations, while effective for single scenes, have proven unsuitable for high-resolution generative models due to their prohibitive memory requirements [6, 36, 47]. Consequently, we adopt the tri-plane representation, which reduces memory demands significantly.
Nevertheless, the tri-plane’s efficacy in representing spatio-temporal data, where the dimensions have different semantic meanings, particularly the temporal axis, requires empirical validation. We conducted a series of overfitting, extrapolation, and interpolation tests to ascertain the viability of the tri-plane representation for video data.
@@
Our research presents three main contributions to the field of video generation: 1) We introduce a tri-plane representation for video data, adept at capturing long-range spatio-temporal dependencies. 2) We develop a generative model design that efficiently handles the creation of extended video sequences. 3) We present a novel optical flow-based motion model that enhances the representation of motion in our generative framework.
These innovations enable us to produce high-quality, photorealistic videos at a resolution of 256x256 pixels for 160 frames at a frame rate of 30 fps. This marks a tenfold improvement in temporal upscaling and a sixteenfold enhancement in spatial resolution, all while requiring three times less computational effort compared to current state-of-the-art models. Moreover, our model supports test-time frame extrapolation and interpolation, which significantly advances the video generation process.
@@ 2 Related work
In the evolving landscape of video generation research, a spectrum of methodologies and frameworks has been proposed to address various facets of the problem. Broadly, the literature can be categorized into three core thematic domains: Content-Guided Video Generation, Computational Efficiency and Training Strategies, and Long-Term Temporal Dynamics. This integrative overview aims to elucidate the seminal contributions in each domain while emphasizing their individual strengths and synergies.
Content-guided video generation
@@
A segment of the literature is devoted to content-guided video generation, which aims to generate videos in adherence to specific content instructions, often furnished through textual prompts or visual descriptors. Among these, the paper on Structure and Content-Guided Video Synthesis with Diffusion Models [8] offers an innovative solution to the intricate challenge of segregating content and structural elements within a video, thereby granting granular control over individual video attributes. Another noteworthy approach, dubbed Phenaki [32], excels in open-domain video generation from textual prompts and stands out for its robust generalization abilities and spatio-temporal consistency. Adding to this category, GODIVA [35] employs a pretrained model to tackle the computational challenges inherent to text-to-video synthesis. This model boasts fine-tuning capabilities and offers zero-shot performance on unseen textual descriptions, fortifying the scope of content-guided video synthesis.
Computational efficiency and training strategies
@@
Efficiency in computational resources and training paradigms constitutes another critical avenue of investigation. For instance, MagicVideo [45] prioritizes computational efficiency by adopting a low-dimensional latent space combined with a 3D U-Net architecture, particularly focusing on the generation of high-resolution videos. StyleGAN-V [28] presents a paradigm shift by conceptualizing video generation as a continuous-time signal, leveraging sparse training data to enhance computational efficiency. Additionally, Video Diffusion Models [16] extend the architecture proposed for image diffusion to facilitate video generation, aiming for temporal coherence and high-fidelity outputs. This model capitalizes on a joint training regimen utilizing both image and video data, thereby accelerating optimization processes and enhancing video quality.
Long-Term temporal dynamics
@@
Another subset of the research corpus addresses the inherent complexity of modeling long-term temporal dynamics in video sequences. The paper titled “Generating Long Videos of Dynamic Scenes” [4] pioneers in this realm by focusing on the accurate reproduction of object trajectories, viewpoint alterations, and content changes over extended time frames. It innovatively refactors temporal latent variables and undergoes training on lengthier videos to circumvent limitations in contemporary methodologies. MoCoGAN [30], on the other hand, proposes a comprehensive framework that separates video into distinct content and motion components, employing adversarial learning techniques along with specialized image and video discriminators to validate its efficacy.
@@
In summary, our work contributes to this body of research by introducing a tri-plane representation of video data, which efficiently models long-range spatio-temporal correlations. We also propose an efficient generator architecture specifically designed to handle extended video sequences. Moreover, we introduce an optical flow-based motion model to facilitate the creation of unconditional, photorealistic videos at a resolution of $256\times 256$ pixels spanning 256 frames. This design further accommodates test-time frame extrapolation and interpolation capabilities.
@@ 3 Method
RAVEN is an unconditional model for generating videos, featuring a generator $\mathcal{G}$ that outputs a complete 256-frame video $\mathcal{V}$ from a single input noise vector $z$ drawn from a 512-dimensional standard normal distribution, denoted as $z\sim\mathcal{N}(\mathbf{0},I^{512\times 512})$. Formally, we express this as $\mathcal{V}=\mathcal{G}(z)$. The generator $\mathcal{G}$ is equipped with a backbone network responsible for producing tri-plane representation features and includes a flow module to enhance the generator’s capacity to represent motion. We use two separate discriminators wich together with the generator are trained using the adversarial objective.
@@ 3.1 Tri-plane hybrid video representation
Generating high-resolution videos with machine learning models necessitates an efficient data representation due to the impractical computational demand of directly applying 3D convolutions. In this section, we adapt the tri-plane representation from EG3D [6] for video data, a technique initially proposed to efficiently represent 3D objects. A suitable representation of 3D objects for generative modeling faces similar memory complexities as ours. Videos, like 3D objects, can be considered three-dimensional arrays, with videos being arrays of pixels and objects as arrays of material properties (density and radiance properties).
@@
This tri-plane approach is more memory-efficient than a full 3D pixel-based video representation, as it can represent a whole video with only three 2D arrays of features. It is also more memory efficient than other implicit video representations, as it uses a simple decoder network and effectively uses the representation capacity of the feature planes. To confirm the efficacy of this representation for videos, we assess its performance through frame interpolation Tab. 1) and extrapolation Tab. 1) experiments, using SSIM and PSNR as our quality metrics, and comparing it against both fully implicit [2] and dense feature volume representation of equivalent memory usage. Results show that the tri-plane representation, especially when augmented with a flow module, outperforms other forms of representation across various datasets and tasks.
Talking Faces all Motions Fashion Videos UCF 101
@@
Pos. emb 0.14 14.43 0.15 14.34 0.66 15.54 0.67 15.60 0.19 14.58 0.18 14.49
voxels 0.75 26.29 0.74 26.08 0.89 26.62 0.88 26.09 0.53 20.61 0.53 20.31
triplane 0.88 31.19 0.87 30.48 0.95 32.38 0.93 30.10 0.58 20.99 0.57 20.65
triplane+flow 0.92 32.77 0.90 32.08 0.96 32.25 0.95 30.71 0.59 21.33 0.58 20.81
@@
Talking Faces all Motions Fashion Videos UCF 101
@@
Pos. emb 0.17 15.26 0.17 15.19 0.70 17.41 0.70 17.59 0.17 14.29 0.21 14.39
voxels 0.73 25.71 0.71 24.77 0.86 24.87 0.85 23.95 0.46 19.54 0.46 19.57
triplane 0.83 28.49 0.82 27.61 0.91 27.65 0.90 26.40 0.50 19.87 0.50 19.88
triplane+flow 0.86 29.76 0.84 28.32 0.92 28.25 0.90 26.65 0.51 20.25 0.52 20.11
@@ 3.2 Tri-Plane + Flow: Explicit motion modeling
We begin by generating the tri-plane feature grids $F_{tri}$ and a separate global feature plane $F_{G}$, derived from the output of the StyleGAN network. At each point in a $64\times 64\times 64$ grid, we sample features from the tri-plane grid, referred to as motion features. These motion features are then decoded by a two-layer MLP (labeled as ‘Flow decoder’ in Fig. 2, $f_{fd}:\mathbb{R}^{36}\rightarrow\mathbb{R}^{5}$) into, local and global flow vectors and a masking scalar (together represented as ‘Flow frames in Fig. 2’).
We obtain the time-dependent feature frame $F_{r}(t)$ as follows. First, using the global flow field $f^{G}_{t}$ at a given time $t$, we warp the global feature frame $F_{G}$ to obtain the time-dependent global features $F_{r}^{G}(t)=w(F_{G},f^{G}_{t})$. Here $w$ is the forward warping function. Concurrently, the local flow $f^{L}_{t}$ helps adjust the preceding feature frame $F_{r}(t-1)$ of time step $(t-1)$ to obtain the time-dependent local features $F_{r}^{L}(t)=w(F_{r}(t),f^{l}_{t})$. These local and global features are alpha-blended to obtain the time-dependent feature at time $t$ as $F_{r}(t)=m_{t}*F_{r}^{G}(t)+(1-m_{t})*F_{r}^{L}(t)$. The mask $m_{t}$ is obtained, as $m_{t}=f_{fd}(F_{tri},t)$, is the output of the flow decoder $f_{fd}$. Mask $m_{t}$ is an additional output alongside the flow fields. By sequentially performing this process for all time steps, we create a $64\times 64\times 64$ grid of appearance features. We call this grid as the appearance volume.
To determine a pixel’s color, we interpolate a feature from the appearance volume at spatio-temporal location of the pixel and decode it with a network comprising two StyleGAN synthesis blocks (labeled simply as ‘Decoder’ in Fig. 2). The incorporation of flow fields into our model shows marked improvements in frame interpolation (Tab. 1) and extrapolation (2) tasks.

### v2

@@ Abstract
We present a novel unconditional video generative model designed to address long-term spatial and temporal dependencies, with attention to computational and dataset efficiency. To capture long spatio-temporal dependencies, our approach incorporates a hybrid explicit-implicit tri-plane representation inspired by 3D-aware generative frameworks developed for three-dimensional object representation and employs a single latent code to model an entire video clip. Individual video frames are then synthesized from an intermediate tri-plane representation, which itself is derived from the primary latent code. This novel strategy more than halves the computational complexity measured in FLOPs compared to the most efficient state-of-the-art methods. Consequently, our approach facilitates the efficient and temporally coherent generation of videos. Moreover, our joint frame modeling approach, in contrast to autoregressive methods, mitigates the generation of visual artifacts. We further enhance the model’s capabilities by integrating an optical flow-based module within our Generative Adversarial Network (GAN) based generator architecture, thereby compensating for the constraints imposed by a smaller generator size. As a result, our model synthesizes high-fidelity video clips at a resolution of $256\times 256$ pixels, with durations extending to more than $5$ seconds at a frame rate of 30 fps. The efficacy and versatility of our approach are empirically validated through qualitative and quantitative assessments across three different datasets comprising both synthetic and real video clips. We will make our training and inference code public.
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Figure 1: Here are samples of frames generated by our method, trained on three different datasets: Talking Faces, Webvid10M-Flowers, and Fashion videos. We select six consecutive frames spaced about 0.8 seconds apart to emphasize the motion between frames in the generated video. Then, we swap the red color channel between each pair of consecutive frames. Specifically, the red channel of the 0th frame is swapped with the 1st, the 2nd with the 3rd, and the 4th with the 5th. This results in three overlapping frames with visible motion across the color channels. The videos are available in the supplementary material.
1,2 Equal contribution.
@@ 1 Introduction
The field of video generation has made remarkable progress in recent years. This development has especially focused on the generation of high-quality content-guided videos that are both realistic and efficient. This progress, however, has not reached the same heights as image generation due to two primary challenges: the inherent complexity of video data and the intensive computational demand of video processing. We choose a GAN-based approach over more performant diffusion-based ones keeping computational and dataset demands in mind.
Most existing unconditional video generation methods leverage high-performing image generation models. To create videos, they generate one frame at a time autoregressively. This technique, while effective, is widely accepted to often accumulate error over the regression chain. This is also known as exposure bias [27, 11, 25, 1]. Moreover, image generators are often imperfect, even when they are carefully trained with underlying 3D structures in narrow domain [17]. This further complicates building autoregressive models that use pre-trained image generators. Content-guided video generation, on the other hand, relies on additional modalities and is dependent on training with extensive datasets and complex models [46, 12].
@@
Our methodology begins by drawing parallels between the representation of 3D objects in neural networks and the challenges faced by video generative models. A 3D object, when modeled as a Neural Radiance Field (NeRF) [28], is described as a continuous volumetric scene function.
By analogy, a video is a 3D array of color values indexed by two spatial coordinates and one temporal coordinate. Thus, representing a video with a parametric function involves mapping a 3D input (x, y, t) to RGB pixel values, effectively parameterizing a video as a continuous volume of color.
When designing a generative model with a continuous output space, various forms can be employed, such as explicit voxel grids [14, 18] or implicit representations [9, 30], which are prevalent in 3D object modeling. However, these representations, while effective for single scenes, have proven unsuitable for high-resolution generative models due to their prohibitive memory requirements [10, 45, 56]. Consequently, we adopt the tri-plane representation, which reduces memory demands significantly. There have also been studies that explore the effectiveness of multiplane representation of videos such as K-Planes [13], HexPlane [8], and Tensor4D [34]. However, they do not extend them in the context of a generative model.
@@
In summary, our research presents three main contributions to the field of video generation: 1) We introduce a tri-plane representation for video data, adept at capturing long-range spatio-temporal dependencies. 2) We develop a generative model design that efficiently handles the creation of extended video sequences. 3) We present a novel optical flow-based motion model that enhances the representation of motion in our generative framework.
These enable us to produce high-quality, photorealistic videos at a resolution of 256x256 pixels for 160 frames at a frame rate of 30 fps. This marks a tenfold improvement in temporal upscaling and a sixteenfold enhancement in spatial resolution, all while requiring three times less computational effort compared to current state-of-the-art models. Moreover, our model supports test-time frame extrapolation and interpolation, which significantly advances the video generation process.
@@ 2 Related work
In the evolving landscape of video generation research, a spectrum of methodologies and frameworks has been proposed to address various complexities that arise. Broadly, the literature can be categorized into three core thematic domains: Content-Guided Video Generation, Computational Efficiency and Training Strategies, and Long-Term Temporal Dynamics. This overview aims to elucidate the contributions in each domain while emphasizing their strengths and synergies.
@@
Content-guided video generation: Content-guided video generation aims to generate videos in adherence to specific content instructions, often furnished through textual prompts or visual descriptors. Among these, the paper on Structure and Content-Guided Video Synthesis with Diffusion Models [12] offers an innovative solution to the intricate challenge of segregating content and structural elements within a video, thereby granting granular control over individual video attributes. However, it requires a prerecorded video to borrow the content from at test time. Another noteworthy approach, dubbed Phenaki [40], excels in open-domain video generation from textual prompts and stands out for its robust generalization abilities and spatio-temporal consistency. However, it still suffers from the non-rigid deformation of rigid objects and therefore lacks realism. Adding to this category, GODIVA [44] employs a pre-trained model to tackle the computational challenges inherent to text-to-video synthesis. However, it operates with a large number of correlated spatio-temporal tokens and suffers from similar shortcomings as Phenaki. GODIVA boasts fine-tuning capabilities and offers zero-shot performance on unseen textual descriptions. Recently released closed-source approaches like SORA [7] further expand the scope of text-guided video generation, but little details are known.
@@
Computational efficiency and training strategies: Efficiency in computational resources and training paradigms constitutes another critical avenue of investigation. For instance, MagicVideo [54] prioritizes computational efficiency by adopting a low-dimensional latent space combined with a 3D U-Net architecture, particularly focusing on high-resolution videos. StyleGAN-V [37] presents a paradigm shift by conceptualizing video generation as a continuous-time signal, leveraging sparse training data to enhance computational efficiency. Additionally, Video Diffusion Models [21] extend the architecture proposed for image diffusion to facilitate video generation, aiming for temporal coherence and high-fidelity outputs. This model capitalizes on a joint training regimen utilizing both image and video data, thereby accelerating optimization processes and enhancing video quality. Albeit more efficient than competitors, these methods typically generate a video one frame at a time and the diffusion-based ones require many denoising steps per frame, making them computationally heavy.
@@
Long-Term temporal dynamics: Another subset of the research corpus addresses the inherent complexity of modeling long-term temporal dynamics in video sequences. The paper titled “Generating Long Videos of Dynamic Scenes” [6] pioneers in this realm by focusing on the accurate reproduction of object trajectories, viewpoint alterations, and content changes over extended time frames. It innovatively refactors temporal latent variables and undergoes training on lengthier videos to circumvent limitations in contemporary methodologies. MoCoGAN [38], on the other hand, proposes a comprehensive framework that separates video into distinct content and motion components, employing adversarial learning techniques along with specialized image and video discriminators to validate its efficacy. The methods discussed above generally follow the setting where temporal correlation among frames comes from the inductive bias that the latent codes from which individual frames are generated are correlated. This however suffers from two major drawbacks. The first is that of computational complexity. This is because the whole model has to run one time for every frame. Second, since every frame is related to its neighbors only via the latent code, it has to account for both fine-grained temporal consistency such as simple pixel-wise translation, and at the same time object-level consistency such as articulated movement or rotation. Since these consistency constraints are not spread across different resolutions of intermediate features, we see non-rigid deformations and unrealistic artifacts sip into the results.
@@
In summary, our work addresses the aforementioned problems by introducing a tri-plane representation of video data, which efficiently models long-range spatiotemporal correlations. We also propose an efficient generator architecture specifically designed to handle extended video sequences. Moreover, we introduce an optical flow-based motion model to facilitate the creation of unconditional, photorealistic videos at a resolution of $256\times 256$ pixels spanning $160$ frames. This design further accommodates test-time frame extrapolation and interpolation capabilities. Finally, since a whole video clip is generated in one shot, the intra-frame correlation is enforced automatically.
@@ 3 Method
RAVEN is an unconditional model for generating videos, featuring a generator $\mathcal{G}$ that outputs a complete 160-frame video clip $\mathcal{V}$ from a single input noise vector $z$ drawn from a 512-dimensional standard normal distribution, denoted as $z\sim\mathcal{N}(\mathbf{0},I^{512\times 512})$. Formally, we express this as $\mathcal{V}=\mathcal{G}(z)$. The generator $\mathcal{G}$ is equipped with a backbone network responsible for producing tri-plane representation features and includes a flow module to enhance the generator’s capacity to represent motion. We use two separate discriminators which together with the generator are trained using the adversarial objective.
@@ 3.1 Tri-plane hybrid video representation
Generating high-resolution videos necessitates an efficient data representation due to the impractical computational demand of direct application of 3D convolutions. In this section, we adapt the tri-plane representation from EG3D [10] for video data, a technique initially proposed to efficiently represent 3D objects. A suitable representation of 3D objects for generative modeling faces similar memory complexities as ours. Videos, like 3D objects, can be considered three-dimensional arrays, with videos being arrays of pixels and objects as arrays of material properties (density and radiance properties).
@@
This tri-plane approach is more memory-efficient than a full 3D voxel-based video representation, as it can represent a whole video with only three 2D arrays of features. It is also more memory efficient than other implicit video representations, as it uses a simple decoder network and effectively uses the representation capacity of the feature planes. To confirm the efficacy of this representation for videos, we assess its performance through frame interpolation (Tab. 1) and extrapolation (Tab. 2) experiments, using SSIM and PSNR as our quality metrics, and comparing it against both fully implicit [3] and dense feature volume representation of equivalent memory usage. Results show that the tri-plane representation, especially when augmented with a flow module, outperforms other forms of representation across various datasets and tasks.
Talking Faces all Motions Fashion Videos WebVid10M-Flowers
@@
Pos. emb 0.14 14.43 0.15 14.34 0.66 15.54 0.67 15.60 0.21 15.56 0.20 15.46
voxels 0.75 26.29 0.74 26.08 0.89 26.62 0.88 26.09 0.59 22.41 0.59 22.48
triplane 0.88 31.19 0.87 30.48 0.95 32.38 0.93 30.10 0.71 24.98 0.71 24.84
triplane+flow 0.92 32.77 0.90 32.08 0.96 32.25 0.95 30.71 0.71 25.33 0.70 24.94
@@
Talking Faces all Motions Fashion Videos WebVid10M-Flowers
@@
Pos. emb 0.17 15.26 0.17 15.19 0.70 17.41 0.70 17.59 0.29 17.51 0.29 17.4
voxels 0.73 25.71 0.71 24.77 0.86 24.87 0.85 23.95 0.49 21.72 0.49 21.82
triplane 0.83 28.49 0.82 27.61 0.91 27.65 0.90 26.40 0.58 23.49 0.57 23.40
triplane+flow 0.86 29.76 0.84 28.32 0.92 28.25 0.90 26.65 0.57 23.57 0.57 23.53
@@ 3.2 Tri-Plane + Flow: Explicit motion modeling
We begin by generating the tri-plane feature grids $F_{tri}$ and a separate global feature plane $F_{G}$, derived from the output of the backbone StyleGAN network. At each point in a $64\times 64\times 64$ grid, we sample features from the tri-plane grid, referred to as motion features. These motion features are then decoded by a two-layer MLP (labeled as ‘Flow decoder’ in Fig. 2, $f_{fd}:\mathbb{R}^{36}\rightarrow\mathbb{R}^{5}$) into, local and global flow vectors and a masking scalar (together represented as ‘Flow frames in Fig. 2’).
We obtain the time-dependent feature frame $F_{r}(t)$ as follows. First, using the global flow field $f^{G}_{t}$ at a given time $t$, we warp the global feature frame $F_{G}$ to obtain the time-dependent global features $F_{r}^{G}(t)=w(F_{G},f^{G}_{t})$. Here $w$ is the forward warping function. Concurrently, the local flow $f^{L}_{t}$ helps adjust the preceding feature frame $F_{r}(t-1)$ of time step $(t-1)$ to obtain the time-dependent local features $F_{r}^{L}(t)=w(F_{r}(t),f^{l}_{t})$. These local and global features are alpha-blended to obtain the time-dependent feature at time $t$ as $F_{r}(t)=m_{t}*F_{r}^{G}(t)+(1-m_{t})*F_{r}^{L}(t)$. The mask $m_{t}$ is obtained, as $m_{t}=f_{fd}(F_{tri},t)$, is the output of the flow decoder $f_{fd}$. Mask $m_{t}$ is an additional output alongside the flow fields. By sequentially performing this process for all time steps, we create a $64\times 64\times 64$ grid of appearance features. We call this grid as the appearance volume. Somewhat similar to our method LEO [43], also relies on optical flow. However, it does so in the image space. Our method on the other hand performs flow-based warping in the latent space. This approach results in only 64 forward warping steps of a $64\times 64$ feature grid to generate a video with $256$ frames. This, compared to $256$ warping steps on a larger grid required by LEO for the same video clip size, is significantly more computationally efficient. Additionally, by operating in the feature space, our method is able to represent complex 3D motions and address occlusion and background dynamics, which are often challenging for traditional optical flow based generation methods.
To determine a pixel’s color, we interpolate a feature from the appearance volume at the spatio-temporal location of the pixel and decode it with a network comprising two StyleGAN synthesis blocks (labeled simply as ‘Decoder’ in Fig. 2). The incorporation of flow fields into our model shows marked improvements in frame interpolation (Tab. 1) and extrapolation (Tab. 2) tasks.